OpenStudy (babyslapmafro):

Please help. How do you find the partial derivative of arctan(x/y) with respect to x?

hartnn (hartnn):

do you know the meaning of partial derivative ?

OpenStudy (babyslapmafro):

yes

hartnn (hartnn):

then have you tried to find the partial derivative ? where r u stuck ?

hartnn (hartnn):

do you know derivative of inverse tan function ?

OpenStudy (babyslapmafro):

\[\frac{ 1 }{ x ^{2}+1 }\]

hartnn (hartnn):

good! now since 'y' is constant, its like taking derivative of inverse tan (ax) {where 'a' =1/y} so, apply the chain rule (know it?) \([\tan^{-1}{ax}]'= \dfrac{1}{1+(ax)^2}\dfrac{d}{dx}(ax)=....?\)

OpenStudy (babyslapmafro):

\[\frac{ 1 }{ 1+(\frac{ 1 }{ y ^{2} }+\frac{ 2 }{ y }x+x ^{2} )}\]

hartnn (hartnn):

where does the expression 1/y^2 +2y/x +x^2 come from ???

OpenStudy (babyslapmafro):

i replaced a with 1/y

hartnn (hartnn):

right, so, it was not (a+x)^2 , it was (ax)^2 so, the denominator would just be \(1+x^2/y^2\) got this ?

hartnn (hartnn):

what about the numerator ?? d/dx (ax) =...?

OpenStudy (babyslapmafro):

numerator = (1/y)

hartnn (hartnn):

no....in chain rule, we need to take derivative of the inner function (here ax or x/y) which is, as you correctly said, 1/y :)

hartnn (hartnn):

one last step of simplification : \(\dfrac{1/y}{1+x^2/y^2}=....?\) multiply numerator and denominator by y^2 to simplify this :)

OpenStudy (babyslapmafro):

y^2/(y^2+x^2)

hartnn (hartnn):

remember that the numerator was 1/y and y^2/y would just be 'y' so your final answer would look like this : \(\large \dfrac{y}{x^2+y^2}\)

OpenStudy (babyslapmafro):

oh right ya, thanks

hartnn (hartnn):

welcome ^_^